# MuRating: A High Quality Data Selecting Approach to Multilingual Large Language Model Pretraining

### v2

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across a wide range of tasks, and recent studies have consistently emphasized the critical role of high-quality pretraining data in driving these advances brown2020language; rae2021scaling; chowdhery2023palm. To improve data quality, various strategies have been adopted, such as deduplication lee2021deduplicating; abbas2023semdedup, heuristic and rule-based filtering rae2021scaling; penedo2024fineweb, and domain-aware sampling xie2023doremi; shen2023slimpajama. While effective, these methods often rely heavily on manual heuristics and domain expertise, lacking a unified or principled framework for evaluating and selecting pretraining data. Moreover, they are typically applied as pre-defined or post-hoc filters, limiting their adaptability to downstream performance. In response, model-based data selection approaches have emerged, aiming to learn data quality judgments from examples or auxiliary supervision. These methods utilize different model architectures and data selection criteria. For instance, DCLM li2024datacomp trains a FastText classifier joulin2016bag using high-quality samples from OH2.5 and Reddit ELI5 as positive supervision, while treating Common Crawl web data as negatives. Other approaches such as AskLLM sachdeva2024train, QuRater wettig2024qurating, and the FineWeb-Edu lozhkov2024fineweb-edu classifier employ prompt-based evaluation criteria using various LLMs to assess the quality of input samples.
@@
Data selection beyond English remains an important challenge fang-etal-2017-learning; laurenccon2022bigscience. While model-based data selection methods have demonstrated effectiveness in improving training quality, they have been developed almost entirely for English and are not explicitly designed or validated for non-English languages, leaving a critical gap in multilingual data quality assessment. As LLMs are increasingly applied in diverse linguistic contexts, there is a growing need for selection strategies that extend beyond English. A recent attempt Fineweb2-HQ messmer2025enhancing train language-specific raters using benchmark datasets as positive supervision and general pretraining corpora as negatives, following a strategy similar to DCLM li2024datacomp. However, this approach uses benchmark-derived data, posing a risk of test set contamination.
@@
In this work, we introduce MuRating, a two-stage, translation-and-pairwise framework for multilingual data-quality estimation. MuRating begins by aggregating multiple state-of-the-art English raters via majority-vote pairwise comparisons, fitting a Bradley–Terry model bradley1952rank to learn a single, unified quality scorer. Next, it translates scored English document pairs into each of 17 target languages and construct monolingual, cross-lingual, and parallel pairs—projecting original preference labels onto translated comparisons and assigning neutral labels to parallel translations. Here, parallel pairs consist of identical content translated into two different languages, while cross-lingual pairs involve distinct texts written in different languages. This design yields one multilingual evaluator that preserves English-derived quality signals while remaining language-agnostic.
@@
We apply MuRating framework to fine-tune a MuRater model to annotate English and multilingual web documents and select top 10% data to pretrain LLaMA-architecture grattafiori2024llama models of 1.2B and 7B parameters for validation. Compared to strong baselines—uniform sampling with 50% more data, QuRater wettig2024qurating, AskLLM sachdeva2024train, FineWeb2-HQ messmer2025enhancing—our selection yields an average gain of 1 to 3.4 points on twelve English benchmarks and 1.8 points on a diverse multilingual suite. We further assess translation fidelity via human evaluation, examine the impact of cross-lingual and parallel data, and compare different score transfer approaches.
@@ 2 Related Work
Data Selection. Data selection is essential in constructing high-quality pretraining corpora for LLMs and typically falls into three main categories: deduplication, heuristic-based filtering, and LLM-guided quality evaluation. Early-stage deduplication removes exact or near-duplicate documents to minimize redundancy and enhance model generalization lee2021deduplicating. More advanced fuzzy and semantic methods filter syntactically or semantically similar content jiang2022fuzzydedup; abbas2023semdedup, which is crucial at scale to avoid training instability and performance degradation xue2023repeat; rae2021scaling.
@@
Heuristic filtering uses rules or lightweight models to exclude low-quality text, such as short, repetitive, or toxic content laurenccon2022bigscience; weber2024redpajama; penedo2023refinedweb; soldaini2024dolma. While handcrafted heuristics can be effective, they often have limited generalization and inefficiency, prompting the use of simple classifiers, perplexity scores, or importance sampling chowdhery2023palm; touvron2023llama; xie2023data; muennighoff2023scaling; li-etal-2024-quantity; li-etal-2024-superfiltering. However, these approaches may unintentionally favor simplistic or repetitive content, which can diminish the diversity and informativeness of the dataset.
@@
In contrast, LLM-guided quality scoring directly leverages language models to evaluate data along dimensions like factuality and coherence gunasekar2023textbooks; sachdeva2024train; li-etal-2024-selective. Frameworks such as QuRating wettig2024qurating and FineWeb-Edu lozhkov2024fineweb-edu prioritize educational content using multi-criteria assessment , while Dataman peng2025dataman and FIRE xu2025fire extend this to domain-aware or reliability-sensitive filtering. Despite their advancement, recent approaches depend heavily on GPT-style judgments, potentially introducing model-specific biases.
@@
Multilingual Pretraining. Efforts to construct multilingual datasets for multilingual LLM pretraining have followed similar strategies to those used for English, incorporating deduplication and heuristic-based filtering techniques. Prominent corpora such as mC4 xue2020mt5, RedPajama weber2024redpajama, CulturalX nguyen2023culturax, HPTL de2024new, and FineWeb-2 penedo2025fineweb2pipelinescale leverage these methods to scale multilingual resources, ranging from a few dozen to thousands of languages, significantly enhancing cross-lingual performance in multilingual LLM pretraining. To mitigate the issue of data scarcity for low-resource languages, TransWeb-Edu wang2024multilingual addresses this by translating high-quality English data into multiple languages. In addition, EMMA-X guo2023emma introduces an EM-inspired framework that jointly learns cross-lingual semantic alignment and sentence representations from large-scale non-parallel multilingual data, while CLIMB guo2025exploring enhances multilingual capability by dynamically adjusting the data-mix ratio across languages. However, research on model-based data selection for multilingual LLM pretraining remains limited. Recently, an initial approach messmer2025enhancing introduced a model-based selection method to refine the FineWeb-2 dataset by training language-specific classifiers, using multilingual benchmark datasets as positive examples and web corpus data as negative examples. However, this method relies on the availability of high-quality samples from existing multilingual benchmarks, which may risk contaminating downstream evaluation tasks with biased data.
@@ 3.1 Integration of English AutoRaters
To convert these pairwise comparisons into continuous scalar quality scores, we employ a learning framework based on the Bradley-Terry model bradley1952rank. Let $s_{\theta}(t)$ denote the learnable scalar quality score of text $t$, parameterized by $\theta$. We adopt a binary cross-entropy loss function, following the formulation proposed in wettig2024qurating, which is analogous to the reward model training paradigm in Reinforcement Learning from Human Feedback (RLHF) ouyang2022training, but without incorporating user prompts or conditioning on input queries:
@@ 3.2.1 Translation-Based Alignment of Multilingual Preferences
This assumption is based on the premise that translation preserves both the semantic content and the relative quality between text pairs. Prior work QuRating wettig2024qurating highlights that pairwise comparisons offer increased stability when evaluating text quality. In multilingual settings, pointwise scoring—where absolute quality scores are assigned to individual texts—is more susceptible to subtle changes in tone or phrasing introduced during translation, which can compromise the consistency of the supervision signal. In contrast, pairwise supervision is inherently more robust to such translation-induced variations. As long as the relative ranking between the texts remains consistent (i.e., $t_{A}^{en}$ continues to be preferred over $t_{B}^{en}$ after translation), the corresponding translated pair $(t_{A}^{m},t_{B}^{m})$ remains a valid training example. This robustness makes pairwise comparisons a more reliable and effective framework for training quality evaluation models in multilingual contexts.
@@ 3.3.2 Training the Rater Model
We then extend these English pairs into multilingual settings using GPT-4o for translation: 150,000 monolingual pairs, 150,000 cross-lingual pairs, and 75,000 parallel pairs $(t_{A}^{m},t_{A}^{m^{\prime}})$, with language proportions balanced across all target languages. The combined dataset—comprising English, monolingual, cross-lingual, and parallel examples—forms the final MuRater training set. We adopt QuRater’s training setup wettig2024qurating, applying a confidence margin to all but the parallel examples.
We fine-tune an encoder-based model following the BGE-M3 architecture bge-m3, adding a linear head to predict quality ratings. We choose BGE-M3 bge-m3 for its strong multilingual representation ability and lightweight design, which make it well-suited for large-scale multilingual scoring. The resulting rater achieves over 93% accuracy on the validation set and 97% on the training set, demonstrating strong multilingual preference modeling. The effect of translation quality and implementation details, including tokenizer settings and hyperparameters, are provided in Appendix 6.
@@ 4.1 Experimental Setups
Dataset construction. We build on the deduplication and heuristic‐filtering pipelines of FineWeb-2 penedo2025fineweb2pipelinescale to assemble a large web‐crawl corpus. It comprises 1.5 trillion English tokens plus 3 trillion tokens across 17 additional languages (Arabic, Chinese, Dutch, French, German, Indonesian, Italian, Japanese, Korean, Portuguese, Russian, Spanish, Thai, Turkish, Vietnamese, Malay, Tagalog). We then apply MuRater and other baselines to assign quality scores to every document. Although scoring trillions of tokens is compute-intensive, it parallelizes efficiently across GPUs, and batching strategies reduce overhead in practice. Corpus statistics are detailed in Appendix 7.1.
@@
Baselines. For the English experiments, we train and evaluate the model using only English pairwise data, comparing it against several established data-quality raters: QuRater, which selects data based on educational value wettig2024qurating; AskLLM, which follows the prompt design in sachdeva2024train using Flan-T5-XXL chung2024scaling; the FineWeb-Edu Class ifier [^2] https://huggingface.co/HuggingFaceFW/fineweb-edu-classifier, trained on 450K LLaMA3-70B-Instruct  [^3] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct labels to identify educational content; and DCLM [^4] https://huggingface.co/mlfoundations/fasttext-oh-eli5, a fastText classifier trained on high quality dataset to differentiate between informative and low-quality web content.
@@
For the multilingual experiments, we extend the QuRater framework wettig2024qurating to build QuRater-M, which employs GPT-4o to annotate multilingual pairs via relative preference, following the same training pipeline as its English counterpart. We further include datasets from HPTL de2024new and FineWeb-2 penedo2025fineweb2pipelinescale for comprehensive comparison, and evaluate against FineWeb2-HQ messmer2025enhancing. In addition, we analyze the difference between MuRater(E) and MuRater(M): MuRater(M) scores multilingual pairs that have been translated into English, whereas MuRater(E) starts from rated English data and translates it into multilingual pair and cross-lingual pair form. Most multilingual evaluations cover 17 languages plus English, while the experiment involving FineWeb2-HQ is restricted to 13 overlapping languages (Arabic, Chinese, Dutch, French, German, Indonesian, Italian, Japanese, Portuguese, Russian, Spanish, Turkish, and Vietnamese) to ensure fair comparison.
@@
Finally, we include a Uniform baseline for both settings, which randomly samples 50% more data than the other methods, following the setup of QuRating wettig2024qurating. Details are provided in Appendix 7.2.
@@
Training Setup. We train a randomly initialized language model based on the LLaMA architecture grattafiori2024llama for a single epoch over the training corpus, with data presented in a randomly shuffled order. For most experiments, the model comprises 1.2 billion parameters and employs a standard transformer architecture vaswani2017attention augmented with rotary position embeddings (RoPE) su2024roformer. To accommodate the multilingual setting, we extend the tokenizer vocabulary through retraining on the multilingual corpus. Comprehensive architectural and tokenizer details are provided in Appendix 7.3.
@@
Evaluation Benchmarks. We assess the performance of our pretrained models using the lm-evaluation-harness framework eval-harness. For the English-only evaluation, we consider a suite of ten tasks spanning multiple linguistic competencies. These include six reading comprehension benchmarks—ARC-Easy, ARC-Challenge clark2018think, SciQ welbl-etal-2017-crowdsourcing, LogiQA liu2020logiqa, TriviaAQjoshi2017triviaqa and BoolQ clark2019boolq; four commonsense reasoning tasks—HellaSwag zellers2019hellaswag, PIQA bisk2019piqa, OpenBookQA OpenBookQA2018 and WinoGrande sakaguchi2019winogrande; and two World Knowledge tasks—Natural Questions (NQ) kwiatkowski2019natural and MMLU hendrycks2021measuring. For MMLU, we follow alzahrani2024benchmarks and employ the lighteval variant to ensure more consistent and reliable comparisons.
For the multilingual evaluation, we utilize translated versions of several English benchmarks han2025mubenchassessmentmultilingualcapabilities alongside multilingual-native datasets. Similar to English benchmarks, we divide the task into three categories, including reading comprehension, commonsense reasoning, and world knowledge understanding. Reading comprehension is evaluated using translated versions of ARC-Easy and ARC-Challenge clark2018think, StoryCloze mostafazadeh2016corpus, and XNLI conneau2018xnli, which assess contextual understanding and inference. Commonsense reasoning is tested with HellaSwag zellers2019hellaswag, XCOPA ponti2020xcopa, and XWinograd tikhonov2021s, focusing on event causality, semantic plausibility, and everyday reasoning across languages. World Knowledge evaluation includes MMLU hendrycks2021measuring, BMLAMA qi-etal-2023-cross, and FLORES goyal2022flores, which examine factual knowledge, translation quality, and multilingual alignment.
To strengthen the evaluation of language-specific knowledge, we also incorporate localized MMLU variants—CMMLU li2024cmmlu, VMLU [^5] https://vmlu.ai/, IndoMMLU koto-etal-2023-indommlu, JMMLU [^6] https://huggingface.co/datasets/nlp-waseda/JMMLU, and AMMLU [^7] https://huggingface.co/datasets/Hennara/ammlu—to construct a region-specific multilingual extension of MMLU, denoted as MMLU_L and add to the world knowledge category. Together, these benchmarks provide a comprehensive evaluation of cross-lingual comprehension, commonsense inference, and knowledge-grounded reasoning, enabling a holistic assessment of multilingual LLM performance. Detailed information of benchmark statistics and language coverage is provided in Appendix 8.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across a wide range of tasks, and recent studies have consistently emphasized the critical role of high-quality pretraining data in driving these advances Brown et al. (2020); Rae et al. (2021); Chowdhery et al. (2023). To improve data quality, various strategies have been adopted, such as deduplication Lee et al. (2021); Abbas et al. (2023), heuristic and rule-based filtering Rae et al. (2021); Penedo et al. (2024), and domain-aware sampling Xie et al. (2023a); Shen et al. (2023). While effective, these methods often rely heavily on manual heuristics and domain expertise, lacking a unified or principled framework for evaluating and selecting pretraining data. Moreover, they are typically applied as pre-defined or post-hoc filters, limiting their adaptability to downstream performance. In response, model-based data selection approaches have emerged, aiming to learn data quality judgments from examples or auxiliary supervision. These methods utilize different model architectures and data selection criteria. For instance, DCLM Li et al. (2024b) trains a FastText classifier Joulin et al. (2016) using high-quality samples from OH2.5 and Reddit ELI5 as positive supervision, while treating Common Crawl web data as negatives. Other approaches such as AskLLM Sachdeva et al. (2024), QuRater Wettig et al. (2024), and the FineWeb-Edu Lozhkov et al. (2024) classifier employ prompt-based evaluation criteria using various LLMs to assess the quality of input samples.
@@
Data selection beyond English remains an important challenge Fang et al. (2017); Laurençon et al. (2022). While model-based data selection methods have demonstrated effectiveness in improving training quality, they have been developed almost entirely for English and are not explicitly designed or validated for non-English languages, leaving a critical gap in multilingual data quality assessment. As LLMs are increasingly applied in diverse linguistic contexts, there is a growing need for selection strategies that extend beyond English. A recent attempt Fineweb2-HQ Messmer et al. (2025) train language-specific raters using benchmark datasets as positive supervision and general pretraining corpora as negatives, following a strategy similar to DCLM Li et al. (2024b). However, this approach uses benchmark-derived data, posing a risk of test set contamination.
@@
In this work, we introduce MuRating, a two-stage, translation-and-pairwise framework for multilingual data-quality estimation. MuRating begins by aggregating multiple state-of-the-art English raters via majority-vote pairwise comparisons, fitting a Bradley–Terry model Bradley and Terry (1952) to learn a single, unified quality scorer. Next, it translates scored English document pairs into each of 17 target languages and construct monolingual, cross-lingual, and parallel pairs—projecting original preference labels onto translated comparisons and assigning neutral labels to parallel translations. Here, parallel pairs consist of identical content translated into two different languages, while cross-lingual pairs involve distinct texts written in different languages. This design yields one multilingual evaluator that preserves English-derived quality signals while remaining language-agnostic.
@@
We apply MuRating framework to fine-tune a MuRater model to annotate English and multilingual web documents and select top 10% data to pretrain LLaMA-architecture Grattafiori et al. (2024) models of 1.2B and 7B parameters for validation. Compared to strong baselines—uniform sampling with 50% more data, QuRater Wettig et al. (2024), AskLLM Sachdeva et al. (2024), FineWeb2-HQ Messmer et al. (2025)—our selection yields an average gain of 1 to 3.4 points on twelve English benchmarks and 1.8 points on a diverse multilingual suite. We further assess translation fidelity via human evaluation, examine the impact of cross-lingual and parallel data, and compare different score transfer approaches.
@@ 2 Related Work
Data Selection. Data selection is essential in constructing high-quality pretraining corpora for LLMs and typically falls into three main categories: deduplication, heuristic-based filtering, and LLM-guided quality evaluation. Early-stage deduplication removes exact or near-duplicate documents to minimize redundancy and enhance model generalization Lee et al. (2021). More advanced fuzzy and semantic methods filter syntactically or semantically similar content Jiang et al. (2022); Abbas et al. (2023), which is crucial at scale to avoid training instability and performance degradation Xue et al. (2023); Rae et al. (2021).
@@
Heuristic filtering uses rules or lightweight models to exclude low-quality text, such as short, repetitive, or toxic content Laurençon et al. (2022); Weber et al. (2024); Penedo et al. (2023); Soldaini et al. (2024). While handcrafted heuristics can be effective, they often have limited generalization and inefficiency, prompting the use of simple classifiers, perplexity scores, or importance sampling Chowdhery et al. (2023); Touvron et al. (2023); Xie et al. (2023b); Muennighoff et al. (2023); Li et al. (2024e, d). However, these approaches may unintentionally favor simplistic or repetitive content, which can diminish the diversity and informativeness of the dataset.
@@
In contrast, LLM-guided quality scoring directly leverages language models to evaluate data along dimensions like factuality and coherence Gunasekar et al. (2023); Sachdeva et al. (2024); Li et al. (2024c). Frameworks such as QuRating Wettig et al. (2024) and FineWeb-Edu Lozhkov et al. (2024) prioritize educational content using multi-criteria assessment , while Dataman Peng et al. (2025) and FIRE Xu et al. (2025) extend this to domain-aware or reliability-sensitive filtering. Despite their advancement, recent approaches depend heavily on GPT-style judgments, potentially introducing model-specific biases.
@@
Multilingual Pretraining. Efforts to construct multilingual datasets for multilingual LLM pretraining have followed similar strategies to those used for English, incorporating deduplication and heuristic-based filtering techniques. Prominent corpora such as mC4 Xue et al. (2020), RedPajama Weber et al. (2024), CulturalX Nguyen et al. (2023), HPTL De Gibert et al. (2024), and FineWeb-2 Penedo et al. (2025) leverage these methods to scale multilingual resources, ranging from a few dozen to thousands of languages, significantly enhancing cross-lingual performance in multilingual LLM pretraining. To mitigate the issue of data scarcity for low-resource languages, TransWeb-Edu Wang et al. (2024) addresses this by translating high-quality English data into multiple languages. In addition, EMMA-X Guo et al. (2023) introduces an EM-inspired framework that jointly learns cross-lingual semantic alignment and sentence representations from large-scale non-parallel multilingual data, while CLIMB Guo et al. (2025) enhances multilingual capability by dynamically adjusting the data-mix ratio across languages. However, research on model-based data selection for multilingual LLM pretraining remains limited. Recently, an initial approach Messmer et al. (2025) introduced a model-based selection method to refine the FineWeb-2 dataset by training language-specific classifiers, using multilingual benchmark datasets as positive examples and web corpus data as negative examples. However, this method relies on the availability of high-quality samples from existing multilingual benchmarks, which may risk contaminating downstream evaluation tasks with biased data.
@@ 3.1 Integration of English AutoRaters
To convert these pairwise comparisons into continuous scalar quality scores, we employ a learning framework based on the Bradley-Terry model Bradley and Terry (1952). Let $s_{\theta}(t)$ denote the learnable scalar quality score of text $t$, parameterized by $\theta$. We adopt a binary cross-entropy loss function, following the formulation proposed in Wettig et al. (2024), which is analogous to the reward model training paradigm in Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022), but without incorporating user prompts or conditioning on input queries:
@@ 3.2.1 Translation-Based Alignment of Multilingual Preferences
This assumption is based on the premise that translation preserves both the semantic content and the relative quality between text pairs. Prior work QuRating Wettig et al. (2024) highlights that pairwise comparisons offer increased stability when evaluating text quality. In multilingual settings, pointwise scoring—where absolute quality scores are assigned to individual texts—is more susceptible to subtle changes in tone or phrasing introduced during translation, which can compromise the consistency of the supervision signal. In contrast, pairwise supervision is inherently more robust to such translation-induced variations. As long as the relative ranking between the texts remains consistent (i.e., $t_{A}^{en}$ continues to be preferred over $t_{B}^{en}$ after translation), the corresponding translated pair $(t_{A}^{m},t_{B}^{m})$ remains a valid training example. This robustness makes pairwise comparisons a more reliable and effective framework for training quality evaluation models in multilingual contexts.
@@ 3.3.2 Training the Rater Model
We then extend these English pairs into multilingual settings using GPT-4o for translation: 150,000 monolingual pairs, 150,000 cross-lingual pairs, and 75,000 parallel pairs $(t_{A}^{m},t_{A}^{m^{\prime}})$, with language proportions balanced across all target languages. The combined dataset—comprising English, monolingual, cross-lingual, and parallel examples—forms the final MuRater training set. We adopt QuRater’s training setup Wettig et al. (2024), applying a confidence margin to all but the parallel examples.
We fine-tune an encoder-based model following the BGE-M3 architecture Chen et al. (2024), adding a linear head to predict quality ratings. We choose BGE-M3 Chen et al. (2024) for its strong multilingual representation ability and lightweight design, which make it well-suited for large-scale multilingual scoring. The resulting rater achieves over 93% accuracy on the validation set and 97% on the training set, demonstrating strong multilingual preference modeling. The effect of translation quality and implementation details, including tokenizer settings and hyperparameters, are provided in Appendix 6.
@@ 4.1 Experimental Setups
Dataset construction. We build on the deduplication and heuristic‐filtering pipelines of FineWeb-2 Penedo et al. (2025) to assemble a large web‐crawl corpus. It comprises 1.5 trillion English tokens plus 3 trillion tokens across 17 additional languages (Arabic, Chinese, Dutch, French, German, Indonesian, Italian, Japanese, Korean, Portuguese, Russian, Spanish, Thai, Turkish, Vietnamese, Malay, Tagalog). We then apply MuRater and other baselines to assign quality scores to every document. Although scoring trillions of tokens is compute-intensive, it parallelizes efficiently across GPUs, and batching strategies reduce overhead in practice. Corpus statistics are detailed in Appendix 7.1.
@@
Baselines. For the English experiments, we train and evaluate the model using only English pairwise data, comparing it against several established data-quality raters: QuRater, which selects data based on educational value Wettig et al. (2024); AskLLM, which follows the prompt design in Sachdeva et al. (2024) using Flan-T5-XXL Chung et al. (2024); the FineWeb-Edu Class ifier [^2] https://huggingface.co/HuggingFaceFW/fineweb-edu-classifier, trained on 450K LLaMA3-70B-Instruct  [^3] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct labels to identify educational content; and DCLM [^4] https://huggingface.co/mlfoundations/fasttext-oh-eli5, a fastText classifier trained on high quality dataset to differentiate between informative and low-quality web content.
@@
For the multilingual experiments, we extend the QuRater framework Wettig et al. (2024) to build QuRater-M, which employs GPT-4o to annotate multilingual pairs via relative preference, following the same training pipeline as its English counterpart. We further include datasets from HPTL De Gibert et al. (2024) and FineWeb-2 Penedo et al. (2025) for comprehensive comparison, and evaluate against FineWeb2-HQ Messmer et al. (2025). In addition, we analyze the difference between MuRater(E) and MuRater(M): MuRater(M) scores multilingual pairs that have been translated into English, whereas MuRater(E) starts from rated English data and translates it into multilingual pair and cross-lingual pair form. Most multilingual evaluations cover 17 languages plus English, while the experiment involving FineWeb2-HQ is restricted to 13 overlapping languages (Arabic, Chinese, Dutch, French, German, Indonesian, Italian, Japanese, Portuguese, Russian, Spanish, Turkish, and Vietnamese) to ensure fair comparison.
@@
Finally, we include a Uniform baseline for both settings, which randomly samples 50% more data than the other methods, following the setup of QuRating Wettig et al. (2024). Details are provided in Appendix 7.2.
@@
Training Setup. We train a randomly initialized language model based on the LLaMA architecture Grattafiori et al. (2024) for a single epoch over the training corpus, with data presented in a randomly shuffled order. For most experiments, the model comprises 1.2 billion parameters and employs a standard transformer architecture Vaswani et al. (2017) augmented with rotary position embeddings (RoPE) Su et al. (2024). To accommodate the multilingual setting, we extend the tokenizer vocabulary through retraining on the multilingual corpus. Comprehensive architectural and tokenizer details are provided in Appendix 7.3.
@@
Evaluation Benchmarks. We assess the performance of our pretrained models using the lm-evaluation-harness framework Gao et al. (2024). For the English-only evaluation, we consider a suite of ten tasks spanning multiple linguistic competencies. These include six reading comprehension benchmarks—ARC-Easy, ARC-Challenge Clark et al. (2018), SciQ Welbl et al. (2017), LogiQA Liu et al. (2020), TriviaAQJoshi et al. (2017) and BoolQ Clark et al. (2019); four commonsense reasoning tasks—HellaSwag Zellers et al. (2019), PIQA Bisk et al. (2019), OpenBookQA Mihaylov et al. (2018) and WinoGrande Sakaguchi et al. (2019); and two World Knowledge tasks—Natural Questions (NQ) Kwiatkowski et al. (2019) and MMLU Hendrycks et al. (2021). For MMLU, we follow Alzahrani et al. (2024) and employ the lighteval variant to ensure more consistent and reliable comparisons.
For the multilingual evaluation, we utilize translated versions of several English benchmarks Han et al. (2025) alongside multilingual-native datasets. Similar to English benchmarks, we divide the task into three categories, including reading comprehension, commonsense reasoning, and world knowledge understanding. Reading comprehension is evaluated using translated versions of ARC-Easy and ARC-Challenge Clark et al. (2018), StoryCloze Mostafazadeh et al. (2016), and XNLI Conneau et al. (2018), which assess contextual understanding and inference. Commonsense reasoning is tested with HellaSwag Zellers et al. (2019), XCOPA Ponti et al. (2020), and XWinograd Tikhonov and Ryabinin (2021), focusing on event causality, semantic plausibility, and everyday reasoning across languages. World Knowledge evaluation includes MMLU Hendrycks et al. (2021), BMLAMA Qi et al. (2023), and FLORES Goyal et al. (2022), which examine factual knowledge, translation quality, and multilingual alignment.
To strengthen the evaluation of language-specific knowledge, we also incorporate localized MMLU variants—CMMLU Li et al. (2024a), VMLU [^5] https://vmlu.ai/, IndoMMLU Koto et al. (2023), JMMLU [^6] https://huggingface.co/datasets/nlp-waseda/JMMLU, and AMMLU [^7] https://huggingface.co/datasets/Hennara/ammlu—to construct a region-specific multilingual extension of MMLU, denoted as MMLU_L and add to the world knowledge category. Together, these benchmarks provide a comprehensive evaluation of cross-lingual comprehension, commonsense inference, and knowledge-grounded reasoning, enabling a holistic assessment of multilingual LLM performance. Detailed information of benchmark statistics and language coverage is provided in Appendix 8.